OpenStudy (anonymous):

Group Theory Q. if G1 is a cyclic group with generator a, prove that G2 is also a cyclic group, with generator f(a). I know I need to show f(ab)=f(a)f(b). but im not sure how to do so given that its general....a cyclic group could be modulo, group of symmetries, etc. Where do I start?

OpenStudy (anonymous):

there must be something missing from this problem

OpenStudy (anonymous):

what is \(G_2\) in relation to \(G_1\)

OpenStudy (anonymous):

oh...yes let f: G1->G2 be an isomorphism

OpenStudy (anonymous):

sorry about that omission

OpenStudy (anonymous):

oh that is rather a lot of information it means they are the same group, only the elements are called different things

OpenStudy (anonymous):

i see. so I just need to show an element from G1 and an element from G2?

OpenStudy (anonymous):

write down exactly what " \(a\) generates the cyclic group \(G_1\)" means

OpenStudy (anonymous):

in G1, <a> = {a1,a2,...an} so f(an)=bn, where bn is some element in g2?? i'm not confindent in the notation

OpenStudy (anonymous):

\(a\) generates the cyclic group \(G_1\) means any element of \(G_1\) is \(a^k\) for some \(k\) i.e. if \(g\in G_1\) then there is a \(k\) such that \(g=a^k\)

OpenStudy (anonymous):

and the notation is \(G_1=<a>\) yes

OpenStudy (anonymous):

oooohhh. yaaa. i forgot that notation for cyclic groups. thank you.

OpenStudy (anonymous):

now what exactly do you want to show?

OpenStudy (anonymous):

if G1 is a cyclic group with generator a, i need to prove that G2 is also a cyclic group, with generator f(a)

OpenStudy (anonymous):

i mean "exactly" that is what most of the work for the beginning group theory problems amount to write down precisely what you want to show, then it will probably be obvious how to get it

OpenStudy (anonymous):

i think it got it. thank you so much. its so easy to not know where to start with these problems, as it is my first upper level math course.

OpenStudy (anonymous):

maybe it is not clear what i mean forget about how to show it for a second what would " G2 is also a cyclic group, with generator f(a)" mean precisely?

OpenStudy (anonymous):

that there is an element in G2, say a, s.t. <a>={a1, a2,...an} so i need to show if y is an element of G2, then y=f(a)=b^k=a^-1??

OpenStudy (anonymous):

it is the same as what i wrote above for \(G_1\) but replacing \(a\) with \(f(a)\)

OpenStudy (anonymous):

i.e. if \(g\in G_2\) then there is an integer \(k\) with \(g=(f(a))^k\)

OpenStudy (anonymous):

that will make \(f(a)\) a generator of \(G_2\)

OpenStudy (anonymous):

its just simply substituting a for f(a)?

OpenStudy (anonymous):

f(a) for a?

OpenStudy (anonymous):

right, that is what you need to show

OpenStudy (anonymous):

so lets pick an arbitrary element \(g\in G_2\)

OpenStudy (anonymous):

there must be some element \(g'\in G_1\) such that \(f(g')=g\) why?

OpenStudy (anonymous):

because its bijective?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

well onto is enough, but yes, bijective is good too now since \(G_1\) is cyclic, what do we know about that \(g'\in G_1\) ?

OpenStudy (anonymous):

i guess the bijective part allows you to say there is a unique element \(g'\in G_1\) such that \(f(g')=g\)

OpenStudy (anonymous):

in any case, what must be true about \(g'\) ?

OpenStudy (anonymous):

this is where you use the fact that \(G_1\) is cyclic and generated by \(a\)

OpenStudy (anonymous):

for g' in G1, there is f(g')=a

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lets go slow \(g'\in G_1\) right? and \(G_1\) is generated by \(a\) this means that \(g'=a^k\) for some \(k\) right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw \(a\in G_1\) not in \(G_2\) it is \(f(a)\in G_2\)

OpenStudy (anonymous):

i just drew a diagram to keep track

OpenStudy (anonymous):

ok so since \(g'=a^k\) what is another way to write \(f(g')=f(a^k)\) using the fact that \(f\) is a homomorphism?

OpenStudy (anonymous):

i haven't learned homomorphism yet, but i suspect f(a^k)=f(a^k)?

OpenStudy (anonymous):

you must have learned about homomorphisms, because an isomorphism is a homomorphism \(f:G_1\to G_2\) is a homomorphism means for all \(x, y \in G_1\) you have \[f(x\circ y)=f(x)\times f(y)\] where \(\circ\) is the multiplication in \(G_1\) and \(\times \) is the multiplication in \(G_2\) sometimes this is abbreviated as \(f(xy)=f(x)f(y)\) but the two multiplications can be different

OpenStudy (anonymous):

homomorphisms is chapter 14. we are on chapter 9 now...isomorphisms.

OpenStudy (anonymous):

this tells you that \[f(a^k)=\left(f(a)\right)^k\]

OpenStudy (anonymous):

ok well that is one property of a isomorphism as well a homomorphism need not be a bijection

OpenStudy (anonymous):

is it clear what the difference between the left and right had side of \[f(a^k)=\left(f(a)\right)^k\] is?

OpenStudy (zzr0ck3r):

hi @joojoo :)

OpenStudy (anonymous):

yes. the left side is cyclic in G1 and right side is cyclic in G2

OpenStudy (anonymous):

hi zz :) group theory is painful.

OpenStudy (anonymous):

not exactly this \(f(a^k)\) means take \(a\) raise it to the power of \(k\) and then send it via \(f\) to \(G_2\)

OpenStudy (anonymous):

yes, i understand that

OpenStudy (anonymous):

this \(\left(f(a)\right)^k\) means take \(a\), map it via \(f\) to \(G_2\) and raise the result to the power of \(k\)

OpenStudy (anonymous):

which means essentially we are done

OpenStudy (anonymous):

i guess that is what i was trying to say...a^k comes from G1 and the left is what lands in G2

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

i mean the right

OpenStudy (anonymous):

so lets recap and put it all in one package

OpenStudy (anonymous):

thank you, satellite...i suspect i will have to stare at this for a while until it sinks in...its not difficult...but i just trip up on writing it all correctly and keeping my elements straight

OpenStudy (zzr0ck3r):

@joojoo told you this place rules:)

OpenStudy (anonymous):

\(G_1\) is cyclic with generator \(a\) and there is an isomorphism \(f:G_1\to G_2\) you want to prove that \(G_2\) is also cyclic with generator \(f(a)\) what this means specifically is that given any \(g_2\in G_2\) you can write \(g_2=\left (f(a)\right)^k\) for some \(k\)

OpenStudy (anonymous):

yes, yes it does, zz!! isomorphs are not fun right now.

OpenStudy (anonymous):

since \(f\) is an isomorphism, there is a \(g_1\in G_1\) such that \(f(g_1)=g_2\)

OpenStudy (anonymous):

since \(G_1\) is cyclic that means \(g_1=a^k\) for some \(k\)

OpenStudy (anonymous):

and since \(f\) is an isomorphism, \(f(g_1)=f(a^k)=\left(f(a)\right)^k=g_2\) which is exactly what you needed to show

OpenStudy (anonymous):

satellite, got it. i will need to practice writing it out on my own....i suspect several times.

OpenStudy (anonymous):

ah ha. i get it.

OpenStudy (anonymous):

when you do any of these types of problems, the first thing you need to do is write down the definitions exactly, i.e. figure out exactly what each statement means without the verbiage "isomorphism", "cyclic" etc. then write down specifically what you know and specifically what you need to show if you are confused, that is fine, that is why you are given the problems. the "prove" part at the beginning should then be more or less obvious

OpenStudy (anonymous):

thank you for your advice. i spend a good amount of time on problems, and more often, i feel like im dragging thru mud. i guess practice will help.

OpenStudy (anonymous):

yes, it will help and later you will look at a problem like this one and think "oh nothing to it, it is much easier than finding the derivative of \(e^x\sin(x)\)

OpenStudy (anonymous):

i hope so :) i really appreciate your patience and help!!

OpenStudy (anonymous):

yw